OpenStudy (anonymous):

What is the equation of the line, in general form, that passes through the point (-1, -1) and is parallel to the line whose equation is x + y = 3? x-y+ 2 = 0 x+y+ 2 = 0 x+y- 2 = 0

OpenStudy (anonymous):

line given is x + y = 3 in y = mx + c form => y = -x+3 So m1 = -1 for a line perpendicular to this, m = -1/m1 so m = -1/-1 in slope intercept form.. (y - y1) = m(x - x1) x1 = -1, y1 = -1 so equation is => (y -1) = 1(x +1) /1 So y-1=x+1 x-y=1+1 x-y=2 So this is the required equation :)

OpenStudy (anonymous):

So would it be x-y +2

jimthompson5910 (jim_thompson5910):

xoya, the question is asking for a parallel line (not a perpendicular one)

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Do you know what it is

OpenStudy (anonymous):

Your welcome :D

jimthompson5910 (jim_thompson5910):

Anything parallel to x+y = 3 is x+y = C where C is some fixed number x+y = C -1 + (-1) = C .... plug in (-1,-1) -2 = C C = -2 So the equation is x+y = -2

OpenStudy (anonymous):

Cool xD

OpenStudy (anonymous):

So it will be x + y + 2 then ok